Grade 11Mechanics

The acceleration of a body moving with velocity u changes with distance x as a=K^2√x where K is a positive constant . The distance travelled by the body when its velocity becomes 2u

Aesha Meghani
8 Years agoGrade 11

1 Answer

Aesha Meghani
8 Years ago
V^2=u^2+2asTherefore4u^2=u^2+2K^2√x×x3u^2/2K^2=x^3/2Therefore x=(3u/2k)4/3Sddghjkkjjhbvvfffghjjvc fyi kkjgffgjjvfjjfj kg hg